OpenStudy (twizttiez):

What is the direct linear variation equation for the relationship? y varies directly with x and y = 4 when x = 12. A. y = x + 8 B. y = x – 8 C. y = 3x D. 1/3x

OpenStudy (twizttiez):

@Data_LG2

OpenStudy (twizttiez):

I think it's either C or D

OpenStudy (igreen):

y = kx Plug in what we know: 4 = k(12) 4 = 12k Divide 12 to both sides

OpenStudy (astrophysics):

professor please

OpenStudy (igreen):

xD

OpenStudy (twizttiez):

One second

OpenStudy (jhannybean):

direct variation: \(y=kx\) indirect variation: \(y=\dfrac{k}{x}\)

OpenStudy (twizttiez):

Okay one second sorry xD

OpenStudy (igreen):

\(k = \dfrac{4}{12}\) Simplify 4/12

OpenStudy (twizttiez):

0.3333333? Thats what i got

OpenStudy (twizttiez):

or 1/3

OpenStudy (twizttiez):

@iGreen

OpenStudy (twizttiez):

Ohhhh nvm xD it's D